OpenStudy (anonymous):

how to do: cos(2a)=-sin(a) with 0

myininaya (myininaya):

i believe you mean pi and not pie -sin(a)=-cos(pi/2-a) or sin(-a)=cos(pi/2+a) so we have cos(2a)=cos(pi/2+a) \[2a=\frac{\pi}{2}+a =>a=\frac{\pi}{2}\] so that is one value that works

myininaya (myininaya):

there is another way to go about it

myininaya (myininaya):

i see three intersections in my calculator all i can get is just pi/2 so let me think of another way

myininaya (myininaya):

sin(a)=1 when a=pi/2 sin(a)=-1/2 when a=7pi/6 or 11pi/6

myininaya (myininaya):

three values! :)

OpenStudy (anonymous):

very cool!

myininaya (myininaya):

:)